A purse contains $3.65 in quarters and dimes. There are 9 more quarters that dimes. Find the number of quarters and dimes.

Answer by stanbon(75887)

A purse contains $3.65 in quarters and dimes. There are 9 more quarters that dimes. Find the number of quarters and dimes.
------
Equations::
Quantity Eq: q = d + 9
Value Eq:::25q+10d = 365 cents
------------------------------------
Rearrange:
q - d = 9
25q+10d = 365
------
Modify for elimination:::
25q - 25d = 225
25q + 10d = 365
-----------
35d = 140
d = 4 (# of dimes)
-------
Solve for "q":
q = d + 9
q = 4 + 9 = 13 (# of quarters)
==================================
Cheers,
Stan H.
===============